OpenStudy (anonymous):

Can someone explain direct variation to me? I kind of don't understand the vocab; like constant of variation, etc. Thanks! xx

jimthompson5910 (jim_thompson5910):

direct variation is the idea that if one variable goes up, then another variable goes up put another way, as one variable increases, so does another if one variable decreases, then the other variable decreases what's important about direct variation equations is that they are always of the form y = kx where k is the constant of variation ------------------------------------------------------------------- example: the equation y = 2x is a direct variation equation because as x increases, y increases. As x decreases, y decreases

jimthompson5910 (jim_thompson5910):

I like to remember it as "direct as in what goes up directly affects the other to go up as well" contrast this with "inverse variation" where if one goes up, the other goes down (or vice versa)

OpenStudy (anonymous):

So the constant of variation is kind of like the rate it goes up? For every 1k there is 2k- so the rate is 2?

OpenStudy (anonymous):

2x*

jimthompson5910 (jim_thompson5910):

yes exactly, it's basically how fast one increases as you increase the other

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Oh can you give me an example? Like step by step?

jimthompson5910 (jim_thompson5910):

of a direct variation equation?

jimthompson5910 (jim_thompson5910):

or inverse?

OpenStudy (anonymous):

Both?

jimthompson5910 (jim_thompson5910):

examples direct: y = 7x inverse: y = 62/x did you need me to explain anything about them?

OpenStudy (anonymous):

Here's an example from my homework: y varies directly to x. Find the constant of variation and write the equation of direct variation that relates the two variables. y=-6/7 when x=-18/35

jimthompson5910 (jim_thompson5910):

"y varies directly to x." what does that mean in terms of an equation?

jimthompson5910 (jim_thompson5910):

does it mean that y = k/x? or that y = kx?

OpenStudy (anonymous):

y=kx

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so if y = kx and y=-6/7 when x=-18/35, then how can we find k?

OpenStudy (anonymous):

k=y/x?

jimthompson5910 (jim_thompson5910):

how about replacing x and y with the values they give us then solve for k?

jimthompson5910 (jim_thompson5910):

y = kx -18/35 = k*(-6/7) now solve for k

OpenStudy (anonymous):

But how can I divide a fraction with a fraction? I don't get it -__-

OpenStudy (anonymous):

Dividing a fraction: -18/35 divided by -6/7 is the same thing as multiplied by the reciprocal of -6/7, which is -7/6. multiply the numerators and the denominators to multiply the fractions.

OpenStudy (anonymous):

Also, since we have 2 negatives, the answer is positive.

OpenStudy (anonymous):

Can you do that step by step?

OpenStudy (anonymous):

OK.

OpenStudy (anonymous):

\[\frac{ 18 }{ 35 } * \frac{ 7 }{ 6 }\]

OpenStudy (anonymous):

multiply 18 by 7 multiply 35 by 6

OpenStudy (anonymous):

so, 126/210.

OpenStudy (anonymous):

isn't it -7?

OpenStudy (anonymous):

Yes, but we can cancel the negative signs b/c we're multiplying a negative by a negative, which will yield a positive no matter what.

OpenStudy (anonymous):

Oh. Okay so then what does all of that mean? So 126/210 is k?

OpenStudy (anonymous):

Yes. Simplify 126/210, though. (Divide num. and denom. by 2: 63/105)

OpenStudy (anonymous):

Ugh. Okay, that's so complicated >.< Thanks!

OpenStudy (anonymous):

haha. Yeah, a lot of math is complicated.

jimthompson5910 (jim_thompson5910):

you need to reduce further 63/105 can be reduced further

OpenStudy (anonymous):

by 3

jimthompson5910 (jim_thompson5910):

divide both by 7 to reduce

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

or 7. sorry.

OpenStudy (anonymous):

Can you simplify it completely?

OpenStudy (anonymous):

Also, I have another question about point slope form! What if I get this answer? 2x+1y=14 I can't have fractions, but x has to be one

OpenStudy (anonymous):

you cant have fractions in standard form. point slope form is fine with it.

OpenStudy (anonymous):

What if I have to put it in standard form?

OpenStudy (anonymous):

Then 2x+y=14 is in standard form.

OpenStudy (anonymous):

And no fractions.

OpenStudy (anonymous):

And x is positive. So it's fine.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Oh that's right! I thought for a second that x had to be one but it doesn't, it only has to be positive(:

OpenStudy (anonymous):

Yup. Any other questions?

OpenStudy (anonymous):

Nope(: Gheez you're good

OpenStudy (anonymous):

I'm gonna have another question on literal equations soon, so heads up xD

OpenStudy (anonymous):

OK :)